Problem 6-10

Ivanhoe Inc. owns and operates a number of hardware stores in the New England region. Recently, the company has decided to locate another store in a rapidly growing area of Maryland. The company is trying to decide whether to purchase or lease the building and related facilities.

Purchase: The company can purchase the site, construct the building, and purchase all store fixtures. The cost would be $1,855,000. An immediate down payment of $414,500 is required, and the remaining $1,440,500 would be paid off over 5 years at $361,600 per year (including interest payments made at end of year). The property is expected to have a useful life of 12 years, and then it will be sold for $501,500. As the owner of the property, the company will have the following out-of-pocket expenses each period.

Property taxes (to be paid at the end of each year)

$41,810

Insurance (to be paid at the beginning of each year)

26,960

Other (primarily maintenance which occurs at the end of each year)

16,340

$85,110


Lease: First National Bank has agreed to purchase the site, construct the building, and install the appropriate fixtures for Ivanhoe Inc. if Ivanhoe will lease the completed facility for 12 years. The annual costs for the lease would be $252,470. Ivanhoe would have no responsibility related to the facility over the 12 years. The terms of the lease are that Ivanhoe would be required to make 12 annual payments (the first payment to be made at the time the store opens and then each following year). In addition, a deposit of $92,800 is required when the store is opened. This deposit will be returned at the end of the 12th year, assuming no unusual damage to the building structure or fixtures.

Click here to view factor tables

Compute the present value of lease vs purchase. (Currently, the cost of funds for Ivanhoe Inc. is 9%.) (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)

Lease Purchase
Present value   $

$


Which of the two approaches should Ivanhoe Inc. follow?

Ivanhoe Inc. should ??

leasepurchase

the facilities
Answer #1
Lease Purchase
Present Value ($2,030,382.84) ($2,269,521.34)

Considering above lease option is viable since the net present value of cash outflow is lower.

Note : We have not been provided the time period for construction of building. In first option when company buys his cash outflow starts from the moment land is bought & even during construction time of building. On other hand under lease option the lease payment starts once the shop is available for use, however such information is not provided hence we have assumed no time lag for this.

The amount payable at the start of year will be included in the starting period i.e. zero year for investment of 414,500 & property tax and the amount paid at the end of period will be considered for next year i.e. in 1st year for Property tax and other expenses :

NPV Calculation of Purchase Option :

Year Investment Property Tax Insurance Others Total
0 (414,500.00)         (26,960.00)       (441,460.00)
1 (361,600.00)         (41,810.00)         (26,960.00) (16,340.00)       (446,710.00)
2 (361,600.00)         (41,810.00)         (26,960.00) (16,340.00)       (446,710.00)
3 (361,600.00)         (41,810.00)         (26,960.00) (16,340.00)       (446,710.00)
4 (361,600.00)         (41,810.00)         (26,960.00) (16,340.00)       (446,710.00)
5 (361,600.00)         (41,810.00)         (26,960.00) (16,340.00)       (446,710.00)
6         (41,810.00)         (26,960.00) (16,340.00)         (85,110.00)
7         (41,810.00)         (26,960.00) (16,340.00)         (85,110.00)
8         (41,810.00)         (26,960.00) (16,340.00)         (85,110.00)
9         (41,810.00)         (26,960.00) (16,340.00)         (85,110.00)
10         (41,810.00)         (26,960.00) (16,340.00)         (85,110.00)
11         (41,810.00)         (26,960.00) (16,340.00)         (85,110.00)
12    501,500.00         (41,810.00) (16,340.00)         443,350.00
Total    (2,269,521.34)

NPV Calculation of Lease Option :

Year Investment Deposit Total
0 (252,470.00)         (92,800.00)       (345,270.00)
1 (252,470.00)       (252,470.00)
2 (252,470.00)       (252,470.00)
3 (252,470.00)       (252,470.00)
4 (252,470.00)       (252,470.00)
5 (252,470.00)       (252,470.00)
6 (252,470.00)       (252,470.00)
7 (252,470.00)       (252,470.00)
8 (252,470.00)       (252,470.00)
9 (252,470.00)       (252,470.00)
10 (252,470.00)       (252,470.00)
11 (252,470.00)       (252,470.00)
12           92,800.00           92,800.00
($2,030,382.84)

I am also providing excel worksheet showing formula :

Table for Purchase option with excel formula :

Year Investment Property Tax Insurance Others Total
0 -414500 -26960 =SUM(B5:E5)
1 -361600 -41810 -26960 -16340 =SUM(B6:E6)
=A6+1 -361600 -41810 -26960 -16340 =SUM(B7:E7)
=A7+1 -361600 -41810 -26960 -16340 =SUM(B8:E8)
=A8+1 -361600 -41810 -26960 -16340 =SUM(B9:E9)
=A9+1 -361600 -41810 -26960 -16340 =SUM(B10:E10)
=A10+1 -41810 -26960 -16340 =SUM(B11:E11)
=A11+1 -41810 -26960 -16340 =SUM(B12:E12)
=A12+1 -41810 -26960 -16340 =SUM(B13:E13)
=A13+1 -41810 -26960 -16340 =SUM(B14:E14)
=A14+1 -41810 -26960 -16340 =SUM(B15:E15)
=A15+1 -41810 -26960 -16340 =SUM(B16:E16)
=A16+1 501500 -41810 -16340 =SUM(B17:E17)
Total =NPV(9%,F6:F17)+F5

The table starts from A4 and ends at F18.

Table for lease option with excel formula :

Year Investment Deposit Total
0 -252470 -92800 =SUM(B24:C24)
1 -252470 =SUM(B25:C25)
=A25+1 -252470 =SUM(B26:C26)
=A26+1 -252470 =SUM(B27:C27)
=A27+1 -252470 =SUM(B28:C28)
=A28+1 -252470 =SUM(B29:C29)
=A29+1 -252470 =SUM(B30:C30)
=A30+1 -252470 =SUM(B31:C31)
=A31+1 -252470 =SUM(B32:C32)
=A32+1 -252470 =SUM(B33:C33)
=A33+1 -252470 =SUM(B34:C34)
=A34+1 -252470 =SUM(B35:C35)
=A35+1 92800 =SUM(B36:C36)
Total =NPV(9%,D25:D36)+D24

The table starts from A23 and ends on D37.